# Learning Decision-Focused Uncertainty Sets in Robust Optimization

### v4

@@ Abstract
We propose a data-driven technique to automatically learn the uncertainty sets in robust optimization. Our method reshapes the uncertainty sets by minimizing the expected performance across a parametric family of problems, subject to guaranteeing constraint satisfaction. Our approach is very flexible, and can learn a wide variety of uncertainty sets while preserving tractability. We solve the constrained learning problem using a stochastic augmented Lagrangian method that relies on differentiating the solutions of the robust optimization problems with respect to the parameters of the uncertainty set. Due to the nonsmooth and nonconvex nature of the augmented Lagrangian function, we apply the nonsmooth conservative implicit function theorem to establish convergence to a critical point, which is a feasible solution of the constrained problem under mild assumptions. Using empirical process theory, we show finite-sample probabilistic guarantees of constraint satisfaction for the resulting solutions. Numerical experiments show that our method outperforms traditional approaches in robust and distributionally robust optimization in terms of out-of-sample performance and constraint satisfaction guarantees.
@@ 1 Introduction
Over the past years, robust optimization (RO) has become a widely adopted efficient tool for decision-making under uncertainty. A robust approach first defines an uncertainty set where the uncertain vector lives, and subsequently hedges its decision against any worst-case realization in this set (Ben-Tal and Nemirovski,, 2008; Ben-Tal et al.,, 2009; Bertsimas et al.,, 2011; Bertsimas and den Hertog,, 2022). To attain proper solution quality and practical robustness the choice of uncertainty set is paramount.
@@
Traditionally, uncertainty sets are shaped and sized to guarantee they contain an appropriately large fraction of the probability mass of the unknown distribution generating the samples (i.e., high coverage). The main appeal of this approach is that the uncertainty set is independent of the subsequent optimization problem considered. In this vein, many approaches to designing uncertainty sets assume structural information of the underlying distributions, and rely on these a priori assumptions to build guarantees of constraint satisfaction (Ben-Tal and Nemirovski,, 2000; Bertsimas et al.,, 2011; Bandi and Bertsimas,, 2012). In particular, a large amount of existing literature studies how to bound the probability of constraint satisfaction based on the size and shape of the uncertainty sets (Li et al.,, 2012; Guzman et al.,, 2017; Bertsimas et al.,, 2021), where the aforementioned structural information is exploited, but the optimization problem itself is not leveraged. By decoupling the design of the uncertainty set from the subsequent optimization problem, the resulting theoretical guarantees can often be quite conservative (Roos and den Hertog,, 2020; Gao and Kleywegt,, 2023).
The trend of constructing uncertainty sets separately from the subsequent optimization problem has continued in the current era of data-driven RO, where the explosion in the availability of data has lead to new paradigms where uncertainty sets are designed directly from data (Bertsimas et al.,, 2018). Whether by combining a priori assumptions with the confidence regions of statistical hypothesis tests on the data (Bertsimas et al.,, 2018), or by approximating a high-probability-region using quantile estimation (Hong et al.,, 2021), significant attention has been given to the probability distribution generating the data and less so to the subsequent optimization problem. More recently in distributionally robust optimization (DRO), the constructed ambiguity set of distributions has historically been required to contain the true data-generating distribution with a certain level of confidence (Esfahani and Kuhn,, 2018). Again, the intuitive appeal of this approach is that the design of the ambiguity set can be carried out in isolation of the subsequent optimization problem.
@@
Unfortunately, decoupling the uncertainty set construction from the subsequent optimization problem in a two-step approach may lead to very conservative decisions. In DRO with Wasserstein ambiguity sets, the two-step approach suffers from the curse of dimensionality as their radius must scale with the dimension of the uncertainty (Weed and Bach,, 2019). On the other hand, one can overcome the curse of dimensionality by exploiting Lipschitz conditions on the objective function of the subsequent optimization problem when designing the ambiguity set (Gao,, 2023). The situation in DRO with $f$-divergence ambiguity sets is even more striking. For instance, Kullback-Leibler ambiguity sets around an empirical distribution can structurally not contain any continuous distribution however large its radius is chosen. Consequently, if the data-generating distribution is continuous a two-step approach is not possible here. Nevertheless, the Kullback-Leibler can be shaped to ensure that the solution has optimal statistical properties when one is willing to exploiting the fact that the subsequent optimization problem has a continuous loss function (Van Parys et al.,, 2021). In both cases, significant reduction in conservatism can be achieved by designing the uncertainty set not only based on the distributions generating the data, but also by exploiting structural properties of the subsequent optimization problem.
The drive for even better performance in RO, then, leads us to a new direction: a method of constructing uncertainty sets that deviates from the conventional decoupled approach and explicitly takes into account the optimization problem as well. This idea has been around for multiple years; as stated by (Ben-Tal et al.,, 2009, page 34), the requirement of containing a high probability mass is only a sufficient (and often restrictive and not needed) condition to guarantee high probability of constraint satisfaction. However, standardized methodology to exploit the notion have seldom been developed. The customary restriction, in essence, allows the uncertainty set developed to hold for all classes of optimization problems, which brings about high levels of conservatism. However, the uncertainty set need only provide high solution quality for the desired optimization problems, and in turn need only include worst-case realizations of uncertainty for these instances.
In this work, we propose a new approach to calibrate uncertainty sets for a particular family of parametric optimization instances. This is a common scenario in many applications, including inventory management, where we solve similar optimization problems with varying initial inventory levels (i.e., the parameters), while satisfying uncertain demand with high probability. Our method minimizes an expected objective over a family of parametric problems, while ensuring data-driven probabilistic guarantees enforced as a conditional value-at-risk constraint (Rockafellar and Uryasev,, 2002). We construct the uncertainty set, which is the family-specific region of worst-case realizations, by learning its shape and size automatically. Specifically, we develop a stochastic augmented Lagrangian approach, relying on nonsmooth implicit differentiation techniques (Agrawal et al., 2019b, ; Agrawal et al., 2019a, ; Bolte et al.,, 2021; Bertrand et al.,, 2022) to compute the derivative of the solution of robust optimization problem with respect to the shape and size of the uncertainty sets. By incorporating the problem objective and the data-driven constraint satisfaction into the learning problem, we control the trade-off between performance and robustness. Our approach offers significant performance improvements over traditional techniques constructing uncertainty sets with high probability mass requirements, and is more flexible than data-driven RO methods that only tune the radius of the uncertainty sets (often via grid search), and not its shape.
@@ 1.1 Our contributions
In this work, we present a new approach to automatically learn the uncertainty sets in RO. We obtain high-quality solutions across a family of parametric optimization problems while ensuring probabilistic guarantees of constraint satisfaction. By directly taking into account the quality of the RO solution, the sets found are decision-focused and eliminate the unnecessary conservatism common in traditional data-driven RO methods.
  • We minimize the expectation of the objective function over a family of parametric RO problems, while ensuring an appropriate probability level of constraint satisfaction. The latter is formulated as constraint on the conditional-value-of-risk ($\mathop{\bf CVaR}$) of the problem family’s uncertain constraints.
  • The aforementioned objective and constraints rely on the solutions of the parametric RO problems. The overall formulation is thus a constrained bi-level problem, with the inner level represented by the parametric RO problems.
  • We develop a stochastic augmented Lagrangian algorithm to solve the constrained learning problem to find the optimal parameters of the uncertainty set. Due to the potential nonsmooth, nonconvex, and therefore nondifferentiable nature of the constructed problem, we show path-differentiability of the Lagrangian, and apply nonsmooth implicit differentiation. We show convergence to a critical point under step-size and boundedness conditions, and establish the critical point as a feasible solution of the constrained problem under mild assumptions.
@@ 1.2 Related work
Data-driven robust optimization has gained wide popularity, by constructing uncertainty sets through data-driven approximations of the unknown data-generating distribution. Using hypothesis testing, Bertsimas et al., (2018) pair a priori assumptions on the distribution with different statistical tests, and obtain various uncertainty sets with different shapes, computational properties, and modeling power. Similar approaches construct data-driven sets using quantile estimation (Hong et al.,, 2021), type-$\infty$ Wasserstein distance (Bertsimas et al.,, 2022), intersections of multiple single-measure-based ambiguity sets (Tanoumand et al.,, 2023), and deep learning clustering techniques (Goerigk and Kurtz,, 2023). All these methods rely on a two-step procedure which separates the construction of the uncertainty set from the resulting robust optimization problem, and often enforces a high coverage requirement on the selected uncertainty sets. As noted by Ben-Tal et al., (2009), however, this requirement may be unnecessary. Recently, Gupta, (2019) proves that while convex problems require DRO ambiguity sets with high coverage to guarantee constraint satisfaction, traditional DRO ambiguity sets for concave problems are significantly larger than the asymptotically optimal sets. In order to find non coverage-based and potentially smaller uncertainty sets that offer improved performance for targeted robust optimization problems, we turn to ideas in differentiable optimization. Recent developments in differentiable convex optimization have enabled a popular end-to-end approach of training a predictive model to minimize loss on a downstream optimization task (Donti et al.,, 2017; Demirović et al.,, 2019; Wilder et al., 2019a, ; Wilder et al., 2019b, ; Cameron et al.,, 2022; Elmachtoub and Grigas,, 2022), for which we also point to the survey by Kotary et al., (2021). Our approach adapts this idea, training instead the shape of the uncertainty set on its performance on parametric robust optimization problems. While developing this work, we became aware of a concurrent work with a similar goal of fitting decision-focused uncertainty sets. In their work, Schuurmans and Patrinos, (2023) construct DRO ambiguity sets which constrain only the worst-case distribution along the direction in which the expected cost of an approximate solution increases most rapidly. Our work, however, considers a broader class of optimization problems and uncertainty sets, and can directly tackle joint chance constraints using hyperparameter tuning.
@@
Contextual optimization and end-to-end approaches, beginning from the Smart Predict, then Optimize framework (Elmachtoub and Grigas,, 2022) have been shown to improve performance as opposed to traditional two-stage approaches, where the model is trained separately from the optimization problem (Costa and Iyengar,, 2023).  Costa and Iyengar, (2023) propose an end-to-end distributionally robust approach for portfolio construction, where they integrate a prediction layer that predicts asset returns with data on financial features and historical returns, and use these predictions in a decision layer that solves the robust optimization problem. Concurrent to our work, Chenreddy and Delage, (2024) propose an end-to-end training algorithm to produce contextual uncertainty sets that achieve a certain conditional coverage. This is to decrease the conservatism of their previous two-stage approach, conditional robust optimization, where they learn a contextually adapted uncertainty set known to contain with high probability the realized uncertainty (Chenreddy et al.,, 2022). We, however, adopt a different notion for context parameters, denoting them family parameters that parameterize different instances of the downstream optimization problem, instead of the uncertain parameter. Our learned uncertainty set therefore does not learn a direct mapping from the context parameters to the uncertainty set (as in contextual and conditional RO (Sadana et al.,, 2023)), as it may be computationally expensive and hard to interpret. We do not enforce coverage constraints, which may introduce unnecessary conservatism, and instead freely shape the uncertainty sets based on their performance on the entire parametralized family of robust problems. As pointed out in the recent survey by Sadana et al., (2023), current work in the end-to-end framework seldom focus on uncertainty in the downstream optimization problem, especially in the constraints. Our approach thus fills this gap, where the process of constructing the uncertainty set is linked in an end-to-end fashion to the solution of the inner robust optimization problem. Our formulation is flexible, as it allows for a wide variety of uncertain constraints and uncertainty sets.
@@
Automatic hyperparameter selection.
Implicit differentiation techniques can greatly accelerate tuning the regularization parameters in regression and classification problems (Bertrand et al.,, 2020, 2022; Blanchet et al.,, 2019). However, most literature focuses on tuning hyperparameters of the objective function in regression and classification tasks, such as hyper-parameters for Lasso regularization (Bertrand et al.,, 2022). In this work, we tune various hyperparameters of the uncertainty sets in RO using implicit differentiation techniques.
@@
Value at risk ($\mathbf{VaR}$) and conditional value at risk ($\mathop{\bf CVaR}$) are common risk measures for quantifying costs encountered in the tails of distributions. A $\mathbf{VaR}_{\eta}$ constraint is a chance constraint that measures the maximum cost incurred with respect to a given confidence level $\eta$, but while intuitive, the function is often intractable, and does not quantify costs beyond the found threshold value. Conditional value at risk, on the other hand, measures the expected cost given that the cost exceeds $\mathbf{VaR}$, and is known to have a tractable formulation (Uryasev and Rockafellar,, 2001; Rockafellar and Uryasev,, 2002). Since its development, $\mathop{\bf CVaR}$ optimization has been a common method for controlling risk in portfolio optimization and option hedging problems, and has also been explored in risk-constrained Markov decision problems, where the risk is presented as a constraint on the $\mathop{\bf CVaR}$ of the cumulative cost (Chow et al.,, 2015). The use of $\mathop{\bf CVaR}$ constraints to substitute chance constraints has also been explored in reliability engineering, where the $\mathop{\bf CVaR}$ constraint has been given a different name: buffered failure probability (Rockafellar and Royset,, 2010). A related concept for controlling risk is the buffered probability of exceedance (bPOE) (Mafusalov and Uryasev,, 2018), which is equivalent to 1 minus the inverse of $\mathop{\bf CVaR}$, and has a level set equivalence with $\mathop{\bf CVaR}$. In our work, we constrain risk, which in RO is the probability of constraint violation, by constraining the $\mathop{\bf CVaR}$ of the uncertain constraint functions.

### v5

@@ Abstract
We propose a data-driven technique to automatically learn contextual uncertainty sets in robust optimization, resulting in excellent worst-case and average-case performance while also guaranteeing constraint satisfaction. Our method reshapes the uncertainty sets by minimizing the expected performance across a contextual family of problems, subject to conditional-value-at-risk constraints. Our approach is very flexible, and can learn a wide variety of uncertainty sets while preserving tractability. We solve the constrained learning problem using a stochastic augmented Lagrangian method that relies on differentiating the solutions of the robust optimization problems with respect to the parameters of the uncertainty set. Due to the nonsmooth and nonconvex nature of the augmented Lagrangian function, we apply the nonsmooth conservative implicit function theorem to establish convergence to a critical point, which is a feasible solution of the constrained problem under mild assumptions. Using empirical process theory, we show finite-sample probabilistic guarantees of constraint satisfaction for the resulting solutions. Numerical experiments show that our method outperforms traditional approaches in robust and distributionally robust optimization in terms of out-of-sample performance and constraint satisfaction guarantees.
@@ 1 Introduction
Over the past years, robust optimization (RO) has become a widely adopted efficient tool for decision-making under uncertainty. A robust approach first defines an uncertainty set where the uncertain vector lives, and subsequently hedges its decision against any worst-case realization in this set (Ben-Tal and Nemirovski,, 2008; Ben-Tal et al.,, 2009; Bertsimas et al.,, 2011; Bertsimas and den Hertog,, 2022). To attain proper solution quality and practical robustness, the choice of uncertainty set is paramount.
@@
Traditionally, uncertainty sets are shaped and sized to guarantee they contain an appropriately large fraction of the probability mass of the unknown distribution generating the samples (i.e., high coverage). The main appeal of this approach is that the uncertainty set is independent of the subsequent optimization problem considered, and guarantees the worst-case performance regardless of the problem. In this vein, many approaches to designing uncertainty sets assume structural information of the underlying distributions, and rely on these a priori assumptions to build guarantees of constraint satisfaction (Ben-Tal and Nemirovski,, 2000; Bertsimas et al.,, 2011; Bandi and Bertsimas,, 2012). In particular, a large amount of existing literature studies how to bound the probability of constraint satisfaction based on the size and shape of the uncertainty sets (Li et al.,, 2012; Guzman et al.,, 2017), where the aforementioned structural information is exploited, but the optimization problem itself is not leveraged. Even in recent data-driven RO, whether by approximating a high-probability-region using quantile estimation (Hong et al.,, 2021), or by learning a contextual uncertainty set known to contain with high probability the realized uncertainty (Messoudi et al.,, 2022), significant attention has been given to the probability distribution generating the data and less so to the quality of the resulting decisions. In distributionally robust optimization (DRO) as well, the constructed ambiguity set of distributions has historically been required to contain the true data-generating distribution with a certain level of confidence (Esfahani and Kuhn,, 2018).
@@
Unfortunately, decoupling the uncertainty set construction from the subsequent optimization problem may lead to very conservative decisions (Roos and den Hertog,, 2020; Gao and Kleywegt,, 2023). In DRO with Wasserstein ambiguity sets, the two-step approach suffers from the curse of dimensionality as their radius must scale with the dimension of the uncertainty (Weed and Bach,, 2019). On the other hand, one can overcome the curse of dimensionality by exploiting Lipschitz conditions on the objective function of the subsequent optimization problem when designing the ambiguity set (Gao,, 2023). The situation in DRO with $f$-divergence ambiguity sets is even more striking. For instance, Kullback-Leibler ambiguity sets around an empirical distribution can structurally not contain any continuous distribution however large its radius is chosen. Consequently, if the data-generating distribution is continuous a two-step approach is not possible here. Nevertheless, the Kullback-Leibler can be shaped to ensure that the solution has optimal statistical properties when one is willing to exploiting the fact that the subsequent optimization problem has a continuous loss function (Van Parys et al.,, 2021). In both cases, significant reduction in conservatism can be achieved by designing the uncertainty set not only based on the distributions generating the data, but also by exploiting structural properties of the subsequent optimization problem.
The drive for even better performance in RO, then, leads us to methods of constructing uncertainty sets that deviate from the decoupled approach, and explicitly take into account the optimization problem. This idea has been around for multiple years; as stated by (Ben-Tal et al.,, 2009, page 34), the requirement of containing a high probability mass is only a sufficient (and often restrictive and not needed) condition to guarantee high probability of constraint satisfaction. In this vein, researchers have turned to constructing uncertainty sets that target directly the probability of constraint violation, and which may have low coverage (Bertsimas and Sim,, 2004; Bertsimas et al.,, 2021; Calafiore and Campi,, 2005). However, while these works take into account the constraints of the optimization problem, and reduce conservatism compared to the decoupled approaches, they either do not optimize over uncertainty-set parameters to find the best one in terms of the performance of the resulting decisions, or only does so in a limited capacity, i.e., with very few tunable parameters.
In fact, the performance of the optimization problem may be judged on criteria that differs across applications, which may be optimized by different parametrizations of the uncertainty sets. Traditionally, the goal of RO uncertainty sets is to provide solutions that perform well with respect to the true worst-case realizations. However, in practical applications, the decision-maker would prefer solutions that perform well in typical cases, not only in the worst-case. Recently, a line of work in Pareto efficiency in RO (Iancu and Trichakis,, 2014; Bertsimas et al.,, 2024; Adelhütte et al.,, 2023) identifies solutions that perform well not only in the worst case, but for all scenarios in the uncertainty set. Extending these ideas, we propose to construct uncertainty sets in a decision-focused manner, such that the resultant robust solutions perform well not only in constraint satisfaction, but also in both worst-case and average-case performance.
@@ 1.1 Our contributions
In this work, we propose a new approach to calibrate uncertainty sets for a family of contextual optimization problems, optimizing over a performance metric for the specific problem family. This is a common scenario in many applications, including inventory management, where we solve similar optimization problems with varying prices (i.e., context parameters), while satisfying uncertain demand with high probability. We construct the uncertainty set, which maps from contexts to context-based regions of uncertainty, by learning its contextual shape parameters. By directly taking into account the quality of the RO solutions using the performance metric, the set found is decision-focused and eliminates the unnecessary conservatism common in existing data-driven RO methods. Our specific contributions are as follows:
  • To train the parameters of the uncertainty set, we minimize the expectation of a performance metric over a family of parametric RO problems, while ensuring an appropriate probability level of constraint satisfaction. The latter is enforced as a constraint on the conditional-value-of-risk ($\mathop{\bf CVaR}$) of the problem family’s uncertain constraints, which are reformulated into a single joint maximum-of-concave constraint. By incorporating all problem objectives and the data-driven constraint satisfaction into the training problem, we control the trade-off between performance and robustness explicitly.
  • The aforementioned performance metric and constraints rely on the solutions of the parametric RO problems. The overall training formulation is thus a constrained bi-level problem, with the inner level represented by the parametric RO problems.
  • We develop a stochastic augmented Lagrangian algorithm to solve the constrained learning problem to find the parameters of the uncertainty set. Due to the potential nonsmooth, nonconvex, and thus nondifferentiable nature of the problem, we show path-differentiability of the Lagrangian, and apply nonsmooth implicit differentiation. We show convergence to a critical point under step-size and boundedness conditions, and establish the critical point as a feasible solution of the constrained problem under mild assumptions.
@@
  • Our uncertainty sets are based on a variety of traditional sets with simple structure, such as ellipsoidal and budget sets. Once the uncertainty set parameters are learned, we obtain a highly tractable formulation.
@@ 1.2 Related work
Data-driven robust optimization has gained wide popularity, by constructing uncertainty sets through data-driven approximations of the unknown data-generating distribution. Using approaches such as quantile estimation (Hong et al.,, 2021), bounding the type-$\infty$ Wasserstein distance (Bertsimas et al.,, 2022), constructing intersections of multiple single-measure-based ambiguity sets (Tanoumand et al.,, 2023), and applying deep learning clustering techniques (Goerigk and Kurtz,, 2023), significant attention has been given to modeling the distribution. In fact, the majority of these methods rely on a two-step procedure which separates the construction of the uncertainty set from the resulting robust optimization problem, and often enforces a high coverage requirement on the selected uncertainty sets. As noted by Ben-Tal et al., (2009), however, this requirement may be unnecessary. Recently, Gupta, (2019) proves that while convex problems require DRO ambiguity sets with high coverage to guarantee constraint satisfaction, traditional DRO ambiguity sets for concave problems are significantly larger than the asymptotically optimal sets. To reduce conservatism, Bertsimas et al., (2018) use hypothesis testing to pair a priori assumptions on the distribution with different statistical tests, and obtain various uncertain-constraint-dependent uncertainty sets with different shapes, computational properties, and modeling power. These sets are non coverage-based, and offer improved performance for the targeted constraints. Similarly, Van Parys et al., (2021) use a DRO approach with constraints on out-of-sample disappointment to find least conservative cost-estimators and optimizers for stochastic optimization problems. We go one step further, and construct parametric uncertainty sets to optimize target performance metrics for particular families of contextual RO problems. To this end, we turn to ideas in differentiable optimization. Recent developments in differentiable convex optimization have enabled a popular end-to-end approach of training a predictive model to minimize loss on a downstream optimization task (Donti et al.,, 2017; Demirović et al.,, 2019; Wilder et al., 2019a, ; Wilder et al., 2019b, ; Cameron et al.,, 2022; Elmachtoub and Grigas,, 2022), for which we also point to the surveys by Kotary et al., (2021) and Mandi et al., (2024). Our approach adapts this idea, training instead the shape of the uncertainty set on its performance on parametric RO problems. While developing this work, we became aware of a concurrent work with a similar goal of fitting decision-focused uncertainty sets. In their work, Schuurmans and Patrinos, (2023) construct DRO ambiguity sets which constrain only the worst-case distribution along the direction in which the expected cost of an approximate solution increases most rapidly. Our work, however, considers a broader class of optimization problems and uncertainty sets, and can directly tackle joint chance constraints using hyperparameter tuning.
@@
Contextual optimization and end-to-end approaches, beginning from the Smart Predict, then Optimize framework (Elmachtoub and Grigas,, 2022) have been shown to improve performance as opposed to traditional two-stage approaches, where the model is trained separately from the optimization problem (Costa and Iyengar,, 2023).  Costa and Iyengar, (2023) propose an end-to-end distributionally robust approach for portfolio construction, where they integrate a prediction layer that predicts asset returns with data on financial features and historical returns, and use these predictions in a decision layer that solves the robust optimization problem. Concurrent to our work, Chenreddy and Delage, (2024) propose an end-to-end training algorithm to produce contextual uncertainty sets that achieve a certain conditional coverage, for robust problems with an uncertain objective. They construct penalty functions for the achieved objective value and the conditional coverage of the uncertainty set, and train the set predictors via gradient descent. We, however, allow for not only an uncertain objective, but also for any number of uncertain constraints, and formulate our problem as a constrained optimization problem. We do not enforce coverage constraints, which may introduce unnecessary conservatism, and instead constrain the $\mathop{\bf CVaR}$ of the constraint violations. We also optimize over the average-case performance in addition to the worst-case performance, which results in less conservative decisions when the uncertainty realization is not worst-case. In addition, we provide convergence and finite sample probabilistic guarantees. As pointed out in the recent survey by Sadana et al., (2023), current work in the end-to-end framework seldom focus on uncertainty in the downstream optimization problem, especially in the constraints. Our approach thus fills this gap.
Pareto efficiency in RO.
The concept of Pareto efficiency in RO was introduced by Iancu and Trichakis, (2014) for linear RO problems, where they identify solutions that yield worst-case optimal performance and which are not dominated by any other such solutions in non-worst-case scenarios. The concept has since been extended to general linear RO problems in Euclidean spaces (Adelhütte et al.,, 2023), and to linear Adaptive Robust Optimization problems (Bertsimas et al.,, 2024). The focus of this line of work is to find Pareto-dominating solutions among a set of RO solutions for a fixed uncertainty set. We use a similar idea, where instead of searching among a set of RO solutions, we optimize a contextual uncertainty set so that the unique RO solution implied by the uncertainty set performs better both in the worst-case and on average.
@@
Value at risk ($\mathbf{VaR}$) and conditional value at risk ($\mathop{\bf CVaR}$) are common risk measures for quantifying costs encountered in the tails of distributions. A $\mathbf{VaR}_{\eta}$ constraint is a chance constraint that measures the maximum cost incurred with respect to a given confidence level $\eta$, but while intuitive, the function is often intractable, and does not quantify costs beyond the found threshold value. Conditional value at risk, on the other hand, measures the expected cost given that the cost exceeds $\mathbf{VaR}$, and, as the tightest convex approximation of $\mathbf{VaR}$, is known to have a tractable formulation (Uryasev and Rockafellar,, 2001; Rockafellar and Uryasev,, 2002). Since its development, $\mathop{\bf CVaR}$ optimization has been a common method for controlling risk in portfolio optimization and option hedging problems, and has also been explored in risk-constrained Markov decision problems, where the risk is presented as a constraint on the $\mathop{\bf CVaR}$ of the cumulative cost (Chow et al.,, 2015). In fact, $\mathop{\bf CVaR}$ is a coherent risk metric (Artzner et al.,, 1999), which satisfies certain axioms that give it value beyond approximating the $\mathop{\bf VaR}$ (Sarykalin et al.,, 2014). The use of $\mathop{\bf CVaR}$ constraints to substitute chance constraints has also been explored in reliability engineering, where the $\mathop{\bf CVaR}$ constraint has been given a different name: buffered failure probability (Rockafellar and Royset,, 2010). In our work, we constrain risk, which in RO is the probability of constraint violation, by constraining the $\mathop{\bf CVaR}$ of the uncertain constraint functions.